and Barb are i e 37% ax bracket or ordnary income and he 20% bracket capital gains ignore e 3 8% additional a on investment income or higher come taxpayers.)They have o ne several lock ofstoc or any years. They considering the sale of two blocks of stock. The sale of one block would produce a gain of $13,000, The sale of the other would produce a loss of $20,000 For purposes of this problem, ignore any restrictions on deductions or credits based on AGI, They have no other gains or losses this year. Read the reguirements Requirement a. How much tax will they save if they sell the block of stock that produces a loss? Paul and Barb will save Requirements How much tax will they save if they sell the block of stock that produces a loss? a. b. How much additional tax vill they pay if they sell the block of stock that produces a gain? c. What will be the impact on their taxes if they sell both blocks of stock?
Answer #1

Answer a:

Annual net capital loss that can be deducted against ordinary Income is $3,000 (for married filing jointly). Excess net capital loss can be carried forward.

Hence tax they will save if they sell block of stock that produces loss = $3,000 * 37% = $1,110

Paul and Berb will save $1,110

Answer (b):

If then sell the block of stock that produces a gain:

Long term capital gain = $13,000

Long term capital gain tax rate = 20%

Additional tax they will pay = $13,000 * 20% = $2,600

Additional tax they will pay = $2,600

Answer (c):

If they sell both block of stocks, long term loss can be adjusted against long term capital gain.

Net long term capital loss = $20,000 - $13,000 = $7,000

Out of this net long term capital loss of $7,000, they can deduct $3,000 against ordinary income of the year and balance they can carry forward.

Hence, combined impact on their tax will be reduction in tax liability = $3,000 * 37% = $1,110

Impact on their tax will be saving in tax payment by $1,110